OpenStudy (anonymous):

No one will help me!!!..Please Help Me!!!! Take the square root of (-5v + 14) and set that equal to v... V=?... In the question it says V is a real number... There is a possibility that there is no solution but I think there is.... PLEASE HELP!!

OpenStudy (anonymous):

hey, can you help me?

geerky42 (geerky42):

Well, we can't solve it if we don't know what -5v + 14 or something equal to. Expression isn't enough, we need equation to solve.

OpenStudy (anonymous):

Is there a way to symplify it though?

geerky42 (geerky42):

well, first I am confused with your question, is that question or is it your work you did for question?

OpenStudy (anonymous):

that is the question... It says "Solve for , where v is a real number. (If there is more than one answer, answer them both.).. Then the equation is: The square root of (-5v+14) = v

geerky42 (geerky42):

Oh okay, so you have the equation \(\sqrt{-5v+14}=v\). What to do first?

OpenStudy (anonymous):

thanks for coming back!

OpenStudy (anonymous):

yes

geerky42 (geerky42):

we don't like square root, right? how to get rid of it?

OpenStudy (anonymous):

by squaring both sides right?

geerky42 (geerky42):

yes, squaring both sides, we now get \(-5v+14 = v^2\). Any idea what to do next? this equation look familiar to you that we can solve?

OpenStudy (anonymous):

- 14 from both sides next?

geerky42 (geerky42):

hmm, yeah, but do that to -5v too

geerky42 (geerky42):

+5v from both sides

OpenStudy (anonymous):

so then, 0=v^2-14+5v?

geerky42 (geerky42):

Right, try to arrange it to like this form 0=ax^2+bx+c

OpenStudy (anonymous):

x^2 + 5v + (-14)?

geerky42 (geerky42):

Yes, you can simplify it to 0 = v^2+5v-14. What next?

OpenStudy (anonymous):

make it (v+?)+(v+?).... right?

geerky42 (geerky42):

Yeah

geerky42 (geerky42):

or you can use quadratic formula. either way works

OpenStudy (anonymous):

what is the quadratic formula?

geerky42 (geerky42):

You just go ahead and factor.

geerky42 (geerky42):

make it into (v+_)(v-_)

OpenStudy (anonymous):

the numbers have to add to equal 5 and multiply to equal -14 right?

geerky42 (geerky42):

right

geerky42 (geerky42):

a+b=5, ab=-14

OpenStudy (anonymous):

(v+7)+(v-2)=0

OpenStudy (anonymous):

now what?

geerky42 (geerky42):

(v+7)(v-2)=0 No + sign in between

OpenStudy (anonymous):

ok, but how do we finish solving it?

OpenStudy (anonymous):

without going back to the original equation?

geerky42 (geerky42):

So we know that (v+7)(v-2)=0 So that mean either (v+7) have to be 0 or (v-2) =0 So split it into v+7=0 and v-2=0

geerky42 (geerky42):

Now can you solve it?

OpenStudy (anonymous):

yeah, the answer is -7 and 2 right?

geerky42 (geerky42):

well, kind of, we are not done though. Now we have to test and verify if v=-7 and v=2 is true or not.

OpenStudy (anonymous):

plug them back into the equation for v.... thanks for helping me!!!!

geerky42 (geerky42):

Try to plug in -7 to original equation \(\sqrt{14-5v}=v\). Is it true?

OpenStudy (anonymous):

it equals positive 7

geerky42 (geerky42):

Left side equals +7, right. what about right side?

OpenStudy (anonymous):

when I plug 7 in for v I get +7 when I plug 2 in for v I get +2

OpenStudy (anonymous):

-7 is the left side so they don't match up

OpenStudy (anonymous):

are you still here?